80. Kip and Yale run separate mining companies in the same forest. Both pollute the river flowing through the forest with debris from their work. In the table below, the first row shows the current level of debris that makes its way into the river from their work. The following table set out information that shows how much it would cost each company to reduce its pollution by additional increments of 10 pounds.

Kip

Yale

Current debris in pounds

60

80

Cost of reducing debris by 10 pounds

$5

$2

Cost of reducing debris by a second 10 pounds

$10

$4

Cost of reducing debris by a third 10 pounds

$15

$6

Cost of reducing debris by a fourth 10 pounds

$20

$8

Cost of reducing debris by a fifth 10 pounds

$25

$10

If each mining company is forced to cut its debris in half, the respective cost to Kip and Yale will be

A. $5, $2

B. $10, $4

C. $30, $20

D. $30, $12
Answer #1

C. $30, $20

(Kip have to cut its debris by 60/2 = 30 pounds. So, Kip's cost = 5 + 10 + 15 = $30
Yale have to cut its debris by 80/2 = 40 pounds. So, Yale's cost = 2 + 4 + 6 + 8 = $20)